please help this is due tonight! thank you!

Check my work 2 Requirement 2: The company has just hired a new marketing manager who insists that unit sales can be dramatically increased by dropping the selling price from $8 to $7. The marketing manager would like to use the following projections in the budget: Part 2 of 2 Year 2 Quarter Year 3 Quarter Data Budgeted unit sales Selling price per unit 50,000 65,000 105,000 65,000 85,00 95,000 15 $7 1 Chapter 8: Applying Excel 2 3 Data Year 3 Quarter 5 Budgeted unit sales 6 7Selling price per unit 8 Accounts 9 Sales collected in the quarter sales are made 10 Sales collected in the quarter after sales are made 11 Desired ending finished goods inventory is 50,000 65,000 105,000 65,000 85,000 95,000 7 per unit receivable, beginning balance $ 65,000 75% 25% 30%) of the budgeted unit sales of the next quarter Graw Hill KPre 012 Next >
Required information 12Finished goods inventory, beginning 13Raw materials required to produce one unit 14Desired ending inventory of raw materials is 15Raw materials inventory, beginning 16Raw material costs 17Raw materials purchases are paid 18 and 19 Accounts payable for raw materials, beginning balance 81,500 12,000 units 5 pounds 10%) of the next quarter's production needs 23,000 pounds 0.80 per pound 60%) in the quarter the purchases are made 40% in the quarter following purchase a. What are the total expected cash collections for the year under this revised budget? ted cash collections for the year
Required information b. What is the total required production for the year under this revised budget? Total required production for the year C. What is the total cost of raw materials to be purchased for the year under this revised budget? Total cost of raw materials to be purchased for the year
Answer #1
Finished Goods Produced
Sales 50000 65000 105000 65000
Closing 19500 31500 19500 25500
Less : - Opening 12000 19500 31500 19500
Production 57500 77000 93000 71000 298500
Raw Material Purchased
Production 287500 385000 465000 355000
Closing 38500 46500 35500 44000
Less : - Opening 23000 38500 46500 35500
303000 393000 454000 363500 1513500
* Rate 0.8 0.8 0.8 0.8
Purchases 242400 314400 363200 290800 1210800
Payable Paid
Opening 81500
Current * 60% 145440 188640 217920 174480
Previous * 40% 81500 96960 125760 145280
Total Payment 226940 285600 343680 319760 1175980

1. Cash Collections would be :-

Opening Accounts Receivable + Sales collection of starting 3 quarters + 75 % of 4th quarter sales

= 65000$ + (50000 + 65000 + 105000) * 7 + (65000 * 7 * 75 %)

= 65000 $ + 1,540,000$ + 341,250$

= 1,946,250 $

2.)   Production would be :-

Sales + Closing - Opening

=(50,000 + 65000 + 105,000+ 65000) + (85000*30%) -12,000

= 285,000 + 25,500 - 12,000

= 298,500 $ ANS

3.) Cost of Raw Material To Be Purchased

  Quarter 1 Year 3 Production =Sales + Closing - Opening

= 85000 + (95000 * 30%) - 25500

= 88,000 units

Raw Material Required = 88000 * 5 = 4,40,000 pounds

Raw Material to be Purchased = ((298500*5 + (88000 * 10%*5) - 23000)

= 1492500 + 44000 -23000 = 1,513,500 pounds

Total Cost = 1513500 * .80 = 1,210,800$

4.) Based on the Table attached above Answer = 1,175,980 $

5.) Yes, it is a potential problem as Budgeted production in 3rd quarter is 93000 which is more than the capacity.

Thank You

Hoping for a Positive Response